Overall status: B configuration (dipole installation 26 Apr (after this no L/S-band polarimetry)). Correlator hw resources support up to 55/24 BlBprs (that many sub-bands or use for additional resolution).
Hardware status and updates (Langley, Perley, Rupen)
EA13 back in the array. EA05 Back to array 18 Apr; EA18 out until 23 May.
Notes (Rupen/Sowinski): Ongoing focus on 3-bit samplers (Dhawan); noise test this week. Need board replacement for B108-t-0. (Personnel: Michael may be back this week; Ken likely at the end of the week).
CBE looking good; CM work to fix an issue recovering from stress/pointing files; CRM to release this week. OPT 1.09 testing under webtest is ongoing.
CASA Update (new functionality available/issues/enhancements):Notes (Miriam): Release delayed until the end of the month; stable delayed until today (decision to avoid impacting the EVLA special issue deadline last Friday).
AIPS Update (new functionality available/issues/enhancements): Notes (Eric): BDF2AIPS task integrated into AIPS.
Notes: Action items in progress (alert on exposure calculator regarding wideband imaging expections; review handling of PIs for upcoming proposals; review if we need to limit the number of such projects to restrict the required support, etc).